# Automorphism groups of affine varieties and their Lie algebras

### v1

@@ 8. Rigid affine varieties
For the following conjectures see e.g. [PZ22, Conjectures 1.0.1 and 1.0.2].
@@
A weaker form of this conjecture is as follows.
Conjecture 8.3.
If an affine variety $X$ admits no effective $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$- and $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$-actions, then the group $\operatorname{Aut}^{0}(X)$ is trivial.
In fact, both conjectures above hold in dimension $2$.
@@
The proof exploits the following fact of independent interest.
Theorem 8.9 ([PR22, Theorem 3.1]).
@@
This theorem implies (a) below; see [PR22, Proposition 5.1 and Remark 5.2] for the proof of (b) and [BG23] for (c). Note that (c) confirms Conjecture 8.2 in certain particular settings.
@@
For a rigid affine variety $X$ the following hold.
@@
  • (c)
    If $X$ is toric or a rational normal affine variety with only constant invertible functions, finitely generated divisor class group and admits a torus action of complexity one, then $\operatorname{Aut}^{0}(X)$ is an algebraic torus.
@@
The following examples show that the latter conclusion does not hold any longer for non-toric affine varieties.
@@
  1. (see [Dani89, Theorem 2].) Let $D$ be the Danielewski surface $\{xy-z^{2}+z=0\}$ in ${\mathbb{A}}^{3}_{{\mathbb{C}}}$ and let $X=D\times{\mathbb{A}}^{1}$. Then $\operatorname{Aut}(X)$ contains a sequence of pairwise non-conjugate algebraic $2$-tori.
@@
Remark 8.14.
Recall that the Koras-Russel threefold $X$ given in ${\mathbb{A}}^{4}_{{\mathbb{C}}}$ by equation $x+x^{2}y+z^{2}+t^{3}=0$ is diffeomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$, but non-isomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$, see e.g. [KML97]. It is neither flexible, nor rigid. The group $\operatorname{Aut}(X)$ is infinite dimensional, acts on $X$ with an open orbit and a unique fixed point (the origin). It coincides with the stabilizer of $X$ in $\operatorname{Aut}({\mathbb{A}}^{4}_{{\mathbb{C}}})$. This group is naturally isomorphic to the subgroup of $\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{C}}})$ (where ${\mathbb{A}}^{3}_{{\mathbb{C}}}=\operatorname{Spec}{\mathbb{C}}[x,z,t]$) of all automorphisms which leave invariant the ideals $(x)$ and $(x^{2},x+z^{2}+t^{3})$, see [DMJP10].
@@ 9.1. Groups acting infinitely transitively on ${\mathbb{A}}^{n}$
For any $n\geq 2$ one can find three $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups of $\operatorname{Aut}({\mathbb{A}}^{n})$ which generate a subgroup acting infinitely transitively on ${\mathbb{A}}^{n}$. The same is true for some $n+2$ root subgroups, that is, $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups normalized by the standard $n$-torus of $\operatorname{Aut}({\mathbb{A}}^{n})$.
@@ 9.2. Tame subgroups
Notice that the de Jonquières subgroup ${\rm Jonq}_{n}\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ (see Remark 6.3) is generated by the standard torus ${\mathbb{T}}$ of $\operatorname{GL}(n,{\mathbb{K}})$ and the unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$. In turn, $\mathcal{J}_{n}$ is generated by the triangular root $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups
@@
consists of locally nilpotent derivations. This is a maximal Lie subalgebra in $\operatorname{Vec}({\mathbb{A}}^{n})$ consisting of locally nilpotent derivations, see [Sku21, Theorems 1 and 7]; cf. also [BPZ22, Theorems 1.4 and 1.5] and the references therein for closely related results. For ${\mathbb{K}}={\mathbb{C}}$, ${\rm Jonq}_{n}$ is a Borel subgroup of $\operatorname{Aut}({\mathbb{A}}^{n})$ that is, a maximal (connected) solvable subgroup, see [FuPo18, Corollary 1.2].
@@
Let $k$ be an arbitrary field. Then $\operatorname{Aut}({\mathbb{A}}^{2}_{k})=\mathop{\rm Tame}_{2}$. Moreover, $\operatorname{Aut}({\mathbb{A}}^{2}_{k})$ is a free product $\operatorname{Aff}_{2}*_{C}{\rm Jonq}_{2}$ amalgamated over $C=\operatorname{Aff}_{2}\cap{\rm Jonq}_{2}$.
@@
$\Phi\colon(x,y,z)\mapsto(x,y,z)+(\sigma z,y+2\sigma x+\sigma^{2}z,0)\quad\text{where}\quad\sigma=x^{2}-yz.$
@@
Indeed, $\partial(\sigma)=0$ and so, $\Phi^{*}(\sigma)=\sigma$. Since also $\Phi^{*}(z)=z$ we have $\Phi\in\operatorname{Aut}_{{\mathbb{K}}[z]}{\mathbb{K}}[z][x,y]$. Nagata showed that
$\Phi\notin\langle\operatorname{Aff}_{2}({\mathbb{K}}[z]),\,{\rm Jonq}_{2}({\mathbb{K}}[z])\rangle$
and suggested that $\Phi\notin\mathop{\rm Tame}_{3}$ is wild, see [Nag72, Conjecture 3.1]. The following theorem due to Shestakov and Umirbaev confirms Nagata’s Conjecture.
@@
Remark 9.8.
A tame subgroup was defined for some other automorphism groups, for instance,
  • for $\operatorname{Aut}(Q_{3})$ where $Q_{3}$ is a smooth affine quadric threefold realized as the underlined variety of $\operatorname{SL}(3,{\mathbb{C}})$, see [LV13];
  • for $\operatorname{Aut}(A_{n})$ where $A_{n}={\mathbb{K}}\langle x_{1},\ldots,x_{n}\rangle$ is the free associative algebra in $n\geq 2$ variables, see [BKY12].
@@
$(x,y,z)\mapsto(x,y,z)+(z,x,0)\sigma\quad\text{where}\quad\sigma=xz-zy,$
which is wild.
Likewise, consider the quadric cone $S_{3}\subseteq{\mathbb{A}}^{4}$ given by equation $\delta:=x_{1}x_{4}-x_{2}x_{3}=0$. There exists an automorphism $\tau$ of ${\mathbb{A}}^{4}$ which restricts to $S_{3}$ and lifts to the Cox ring ${\mathbb{K}}[x_{1},x_{2},x_{3},x_{4}]$ of $S_{3}$ yielding the Anick automorphism of ${\mathbb{A}}_{4}$:
$(x_{1},x_{2},x_{3},x_{4})\mapsto(x_{1},x_{2}+x_{1}\delta,x_{3},x_{4}+x_{3}\delta).$
@@ 9.4. Stable tameness
A wild automorphism $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is called stably tame if $\varphi\oplus{\rm\operatorname{id}}\in\operatorname{Aut}({\mathbb{A}}^{n+\kappa})$ is tame for some $\kappa\geq 1$. The number $\kappa$ of additional variables will be called an excessive dimension.
@@
By [Spo07] the same is true for the Anick automorphism of $A_{3}$, cf. Remark 9.8. More generally, the following holds.
@@
Let as before $\mathfrak{G}\subseteq\operatorname{Aut}({\mathbb{A}}^{3})$ be the subgroup of automorphisms fixing $z$. Then any element $\varphi\in\mathfrak{G}$ is stably tame with excessive dimension $\kappa=3$.
A similar fact is also known for $\operatorname{Aut}(A_{3})$ with $\kappa=1$, see [BKY12, Theorem 1.1].
@@ 9.6. Amalgamation
By Jung-van der Kulk’s Theorem 9.3, $\operatorname{Aut}({\mathbb{A}}^{2})$ is an amalgam of two subgroups along their intersection. A similar decomposition exists for the group $\operatorname{GL}(2,{\mathbb{K}}[x])$, see [Nag59], and the group $\operatorname{Aut}^{0}(A_{2})$ of augmentation-preserving automorphisms of the free associative algebra $A_{2}$, see [SY98, Theorem 1.1].
It is unknown, however, whether a similar fact holds for $\operatorname{Aut}({\mathbb{A}}^{n})$ when $n\geq 3$. For the tame subgroup $\mathop{\rm Tame}_{3}\subseteq\operatorname{Aut}({\mathbb{A}}_{3})$ we have the following result.
@@
$\mathop{\rm Tame}_{3}$ can be realized as a generalized amalgamated product, specifically, the product of three subgroups amalgamated along pairwise intersections.
See [ANU20, Theorem 1] and the references therein for analogs of the Jung-van der Kulk’s Theorem for the automorphism group of the free Lie algebra in three variables and the subgroup of tame automorphisms. Similar facts are also known to hold for some $\operatorname{Aut}$-homogeneous affine surfaces.

### v2

@@ 8. Rigid affine varieties
Example 8.2.
Conjecturally, a Pham–Brieskorn hypersurface $X\subseteq{\mathbb{A}}^{n+1}$, $n\geq 2$, defined by
$x_{0}^{a_{0}}+x_{1}^{a_{1}}+\cdots+x_{n}^{a_{n}}=0\quad\text{where}\quad 2\leq a_{0}\leq a_{1}\leq\cdots\leq a_{n}$
is rigid if and only if $a_{1}\geq 3$, see [CPPZ21, Conjecture 1.22]. This is indeed the case for $n=2$ [KZ00, Lemma 4] and for $n=3$ [ChDu23, Main theorem]. For example, the Fermat threefold hypersurface
$x_{0}^{3}+x_{1}^{3}+x_{2}^{3}+x_{3}^{3}=0$
in ${\mathbb{A}}^{4}$ is rigid; see [CPPZ21, Theorem 1.13] for a far more general result. See also [EGS23] and references therein for a description of rigid trinomial affine varieties.
For the following conjecture see e.g. [PZ22, Conjectures 1.0.1 and 1.0.2]; cf. also Conjecture 5.9 and Question 5.12.
@@
Conjecture 8.3 is confirmed for rigid toric varieties and varieties with a torus action of complexity one, see [BG23]. Due to the following theorem, this conjecture also holds in dimension $2$.
@@
The proof exploits the following result of independent interest.
Theorem 8.9 ([PR24, Theorem 3.1]).
@@
This theorem implies (a) below; see [PR24, Proposition 5.1 and Remark 5.2] for the proof of (b).
@@
Let $X$ be a rigid affine variety. Then the following hold.
@@
The next theorem confirms Conjecture 8.3 in a particular setting of rigid varieties with a torus action. For statement (a) see [AG17, Example 2.3 and Remark 2.4] and [AKZ12, Sect. 2]; (b) is proven in [BG23, Theorem 6.4]. Recall that the complexity of an algebraic group action is the codimension of general orbits.
Theorem 8.12.
  • (a)
    Let $X$ be a normal toric variety. If $X$ is rigid, then $X\cong{\mathbb{T}}$ is a torus and $\operatorname{Aut}^{0}(X)\cong{\mathbb{T}}$ is the torus acting naturally on itself.
  • (b)
    Let $X$ be a normal rational affine variety with a torus action of complexity one and a finitely generated divisor class group. Suppose that $X$ is rigid and every invertible regular function on $X$ is constant. Then $\operatorname{Aut}^{0}(X)$ is an algebraic torus.
@@
However, for a non-toric affine variety $X$ the group $\operatorname{Aut}(X)$ can contain several pairwise non-conjugate maximal algebraic tori.
@@
  1. (See [Dani89, Theorem 2].) Let $D$ be the Danielewski surface $\{xy-z^{2}+z=0\}$ in ${\mathbb{A}}^{3}_{{\mathbb{C}}}$ and let $X=D\times{\mathbb{A}}^{1}$. Then $\operatorname{Aut}(X)$ contains an infinite sequence of pairwise non-conjugate algebraic $2$-tori and no $3$-torus. Furthermore, let $X$ be the underlying affine variety of $\operatorname{SL}(2,{\mathbb{C}})\times{\mathbb{A}}^{1}_{{\mathbb{C}}}$. Then the number of conjugate classes of maximal tori in $\operatorname{Aut}(X)$ is infinite [Dani89, P. 7].
@@
Remark 8.15 (see [KML97] and [DMJP10]).
Recall that the Koras-Russel threefold $X$ given in ${\mathbb{A}}^{4}_{{\mathbb{C}}}$ by equation $x+x^{2}y+z^{2}+t^{3}=0$ is diffeomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$, but non-isomorphic to ${\mathbb{A}}^{3}_{{\mathbb{C}}}$. It is neither flexible, nor rigid. The group $\operatorname{Aut}(X)$ is infinite dimensional, acts on $X$ with an open orbit and a unique fixed point (the origin). It coincides with the stabilizer of $X$ in $\operatorname{Aut}({\mathbb{A}}^{4}_{{\mathbb{C}}})$. This group is naturally isomorphic to the subgroup of $\operatorname{Aut}({\mathbb{A}}^{3}_{{\mathbb{C}}})$ (where ${\mathbb{A}}^{3}_{{\mathbb{C}}}=\operatorname{Spec}{\mathbb{C}}[x,z,t]$) of all automorphisms which leave invariant the ideals $(x)$ and $(x^{2},x+z^{2}+t^{3})$.
@@ 9.1. Groups acting infinitely transitively on ${\mathbb{A}}^{n}$
Recall that the root subgroups of $\operatorname{Aut}({\mathbb{A}}^{n})$ are the $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups normalized by the standard $n$-torus, see Definition 7.8.
@@
For any $n\geq 2$ one can find three $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$-subgroups (resp., $n+2$ root subgroups) of $\operatorname{Aut}({\mathbb{A}}^{n})$ which generate a subgroup acting infinitely transitively on ${\mathbb{A}}^{n}$.
@@ 9.2. Tame subgroups
Notice that the de Jonquières subgroup ${\rm Jonq}_{n}\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$ (see Remark 6.3) is generated by the standard torus ${\mathbb{T}}$ of $\operatorname{GL}(n,{\mathbb{K}})$ and the unipotent radical $\mathcal{J}_{n}=R_{\mathrm{u}}({\rm Jonq}_{n})$. In turn, $\mathcal{J}_{n}$ is generated by the triangular root subgroups
@@
consists of locally nilpotent derivations. This is a maximal Lie subalgebra in $\operatorname{Vec}({\mathbb{A}}^{n})$ consisting of locally nilpotent derivations, see [Sku21, Theorems 1 and 7]; cf. also [BPZ24, Theorems 1.4 and 1.5] and the references therein for closely related results. For ${\mathbb{K}}={\mathbb{C}}$, ${\rm Jonq}_{n}$ is a Borel subgroup of $\operatorname{Aut}({\mathbb{A}}_{{\mathbb{C}}}^{n})$, that is, a maximal (connected) solvable subgroup, see [FuPo18, Corollary 1.2].
@@
Let $k$ be an arbitrary field. Then Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{k})=\mathop{\rm Tame}_{2}$. Moreover, Aut($\operatorname{Aut}({\mathbb{A}}^{2}_{k})$ is a free product $\operatorname{Aff}_{2}*_{C}{\rm Jonq}_{2}$ amalgamated over $C=\operatorname{Aff}_{2}\cap{\rm Jonq}_{2}$.
@@
$\Phi\colon(x,y,z)\mapsto(x+\sigma z,y+2\sigma x+\sigma^{2}z,z)\quad\text{where}\quad\sigma=x^{2}-yz.$
@@
Indeed, $\partial(\sigma)=0$ and so, $\Phi^{*}(\sigma)=\sigma$. Since also $\Phi^{*}(z)=z$ we have $\Phi\in\operatorname{Aut}_{{\mathbb{K}}[z]}{\mathbb{K}}[z][x,y]$. Nagata showed that $\Phi$ is not generated by the affine and triangular automorphisms that fix $z$, and suggested that $\Phi$ is wild, i.e., $\Phi\notin\mathop{\rm Tame}_{3}$, see [Nag72, Conjecture 3.1]. The following theorem due to Shestakov and Umirbaev confirms Nagata’s Conjecture.
@@
Remarks 9.8.
1. For $n=2$ the assertion was established in [Dés06]. Since $\mathop{\rm Tame}_{2}=\operatorname{Aut}({\mathbb{A}}^{2}_{\mathbb{C}})$, see Theorem 9.3, it follows that every automorphism of $\operatorname{Aut}({\mathbb{A}}^{2}_{\mathbb{C}})$ is inner up to a field automorphism. The latter does not hold, in general, for toric affine surfaces. For example, this fails for the quotient of ${\mathbb{A}}^{2}$ by the cyclic group generated by $g\colon(x,y)\mapsto(\xi^{e}x,\xi y)$, where $\xi$ is a $d$th primitive root of unity, $1<e<d$, $(e,d)=1$ and $e^{2}\not\equiv 1\mod d$, see [LRU23, Remark 5.13].
2. Any automorphism of $\operatorname{Aut}({\mathbb{A}}^{n})$ as an ind-group is inner, see [BKYE16]. The same holds for the automorphisms of the subgroup ${\rm SAut}({\mathbb{A}}^{n})$ of volume-preserving automorphisms of ${\mathbb{A}}^{n}$, see [Kra17]. See also [Sta13], [UZ21, Theorem 1.3] and the literature therein for some generalizations of Theorem 9.7.
3. A tame subgroup was defined for some other automorphism groups, for instance,
  • for $\operatorname{Aut}(Q_{3})$ where $Q_{3}$ is a smooth affine quadric threefold realized as the underlying variety of $\operatorname{SL}(2,{\mathbb{C}})$, see [LV13];
  • for $\operatorname{Aut}(A_{n})$ where $A_{n}={\mathbb{K}}\langle x_{1},\ldots,x_{n}\rangle$ is the free associative algebra in $n\geq 2$ variables, see [BKY12b];
  • for $\operatorname{Aut}(X)$, where $X$ is a toric affine variety, see [AG10, Definition 5.3].
@@ 3. A tame subgroup was defined for some other automorphism groups, for instance,
$(x,y,z)\mapsto(x+z\sigma,y+x\sigma,z)\quad\text{where}\quad\sigma=xz-zy,$
which is wild. See also [SZ24] for a wild (but stably tame) automorphism of Anick type of a $3$-generated free Poisson algebra, which induces a tame automorphism of the 3-generated polynomial algebra.
Likewise, consider the quadratic cone $S_{3}\subseteq{\mathbb{A}}^{4}$ given by equation $x_{1}x_{4}-x_{2}x_{3}=0$. There exists an automorphism $\tau$ of ${\mathbb{A}}^{4}$ which restricts to $S_{3}$ and lifts to the Cox ring ${\mathbb{K}}[x_{1},x_{2},x_{3},x_{4}]$ of $S_{3}$ yielding the Anick automorphism of ${\mathbb{A}}^{4}$:
$(x_{1},x_{2},x_{3},x_{4})\mapsto(x_{1},x_{2}+x_{1}\delta,x_{3},x_{4}+x_{3}\delta)\quad\text{where}\quad\delta=x_{1}x_{4}-x_{2}x_{3}.$
@@
The following theorem relates the automorphism group of ${\mathbb{A}}^{n}$ with the automorphism groups of various Lie algebras of vector fields on ${\mathbb{A}}^{n}$. Below $\operatorname{Vec}^{0}({\mathbb{A}}^{n})$ stands for the Lie algebra of vector fields with zero divergence.
Theorem 9.9.
There are isomorphisms
$\operatorname{Aut}({\mathbb{A}}^{n})\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}({\mathbb{A}}^{n}))\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n}))\cong\operatorname{Aut}_{\rm Lie}(\operatorname{Vec}^{0}({\mathbb{A}}^{n}))$
induced by the natural action of the group $\operatorname{Aut}({\mathbb{A}}^{n})$ on the Lie algebra of vector fields $\operatorname{Vec}({\mathbb{A}}^{n})$.
The first isomorphism was obtained in [Rud86, Corollary], [Kul93, Sec. 4] and [Bav17a] along different approaches; in [Kul93] it appeared in relation with the Jacobian Conjecture. The other isomorphisms were established independently in [Bav17b] and [KrRe17]. See [Bav14] for the automorphism group of the Lie algebra of triangular derivations.
See also [CL06, Théorème A] for a description of the lattices in simple real Lie groups that can be realized as subgroups of $\operatorname{Aut}({\mathbb{A}}^{2}_{{\mathbb{C}}})$. Furthermore, see [AGL24] and the references therein for studies on (finite dimensional) Lie algebras of derivations of polynomial rings.
@@ 9.3. Tamizable automorphisms
The locally finite automorphisms of $\operatorname{Aut}({\mathbb{A}}^{n})$ were studied in [FuMa07]. It is known that the subgroup $\langle{\rm LF}_{n}\rangle$ of $\operatorname{Aut}({\mathbb{A}}^{n})$ generated by all locally finite automorphisms is normal, see [Zyg11]. By the Jung-van der Kulk Theorem 9.3 we have $\langle{\rm LF}_{2}\rangle=\operatorname{Aut}({\mathbb{A}}^{2})$. For $n=3$, $\langle{\rm LF}_{3}\rangle$ contains the Nagata automorphism. However, it is unknown whether $\operatorname{Aut}({\mathbb{A}}^{3})=\langle{\rm LF}_{3}\rangle$ ([Zyg11]; cf. also subsection 9.5).
@@ 9.4. Stable tameness
A wild automorphism $\varphi\in\operatorname{Aut}({\mathbb{A}}^{n})$ is called stably tame if $\varphi\times{\rm\operatorname{id}}_{{\mathbb{A}}^{\kappa}}\in\operatorname{Aut}({\mathbb{A}}^{n+\kappa})$ is tame for some $\kappa\geq 1$. The minimal number $\kappa$ of additional variables will be called excessive dimension.
@@
By [Spo07] the same is true for the Anick automorphism of ${\mathbb{A}}^{4}$, cf. Remark 9.8.3. More generally, the following holds.
@@
Let as before $\mathfrak{G}\subseteq\operatorname{Aut}({\mathbb{A}}^{3})$ be the subgroup of automorphisms fixing $z$. Then any element $\varphi\in\mathfrak{G}$ is stably tame with excessive dimension $\kappa\leq 3$.
A similar result is also known for $\operatorname{Aut}(A_{3})$ with $\kappa=1$, see [BKY12b, Theorem 1.1].
@@ 9.6. Amalgamation
By Jung-van der Kulk’s Theorem 9.3, $\operatorname{Aut}({\mathbb{A}}^{2})$ is an amalgam of two subgroups along their intersection. A similar decomposition exists for the group $\operatorname{GL}(2,{\mathbb{K}}[x])$, see [Nag59], and for the group $\operatorname{Aut}^{0}(A_{2})$ of augmentation-preserving automorphisms of the free associative algebra $A_{2}$, see [SY98, Theorem 1.1].
It is unknown, however, whether a similar result holds for $\operatorname{Aut}({\mathbb{A}}^{n})$ when $n\geq 3$. For the tame subgroup $\mathop{\rm Tame}_{3}\subseteq\operatorname{Aut}({\mathbb{A}}_{3})$ we have the following result.
@@
$\mathop{\rm Tame}_{3}$ can be realized as a generalized amalgamated product of three subgroups amalgamated along pairwise intersections.
See [ANU20, Theorem 1] and the references therein for analogues of the Jung-van der Kulk’s Theorem for the automorphism group of the free Lie algebra in three variables and the subgroup of tame automorphisms. Similar results are also known to hold for certain $\operatorname{Aut}$-homogeneous affine surfaces among the so-called Danilov-Gizatullin surfaces.